OpenStudy (anonymous):

Calc question! How do i evaluate the integral of [2dx/(x^2+1)] over -infinity to -2?

OpenStudy (turingtest):

over infinity you say?

OpenStudy (turingtest):

oh you mean\[\int_{-\infty}^{-2}{2dx\over x^2+1}\]

OpenStudy (anonymous):

yeah exactly! i totally forgot how to do improper integrals.

OpenStudy (anonymous):

thanks for replying btw i was freakin out for a bit lol

OpenStudy (turingtest):

we rewrite this as a limit-integral pair

OpenStudy (anonymous):

oops its actually x^2 - 1

OpenStudy (turingtest):

\[\lim_{n\to-\infty}\int_{n}^{-2}{2dx\over x^2-1}\]I suppose we can do partial fractions to do this integral...

OpenStudy (turingtest):

do you have any problems with this as an indefinite integral ?

OpenStudy (anonymous):

its part of that section in the book

OpenStudy (turingtest):

what section? I'm asking if you can do the indefinite integral\[\int{2dx\over x^2-1}\]?

OpenStudy (anonymous):

a little bit i want to say you use trig sub...

OpenStudy (anonymous):

i need help with the entire problem im sorry haha

OpenStudy (turingtest):

I would say partial fractions is the way to go

OpenStudy (turingtest):

brb

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

sorry about that, I hate long phone calls

OpenStudy (turingtest):

anyway we would use partial fractions on this integral, are you familiar with that technique?

OpenStudy (anonymous):

u -sub will work for this one

OpenStudy (turingtest):

really?

OpenStudy (anonymous):

\(u=x^2-1, du=2xdx\) etc

OpenStudy (anonymous):

off hand i would say it diverges

OpenStudy (turingtest):

it's 2dx on the top though

OpenStudy (anonymous):

\[\int\frac{2xdx}{x^2-1}dx=\int\frac{du}{u}=\ln(u)\]

OpenStudy (turingtest):

you are imagining an x :P

OpenStudy (anonymous):

ooooooooooooh silly me you are right!!!

OpenStudy (turingtest):

but yeah this integral still gets funky

OpenStudy (turingtest):

the limit I'm not sure about...

OpenStudy (anonymous):

no i think you may be able to do it. \[\int \frac{dx}{x-1}-\int \frac{dx}{x+1}\] \[\ln(x-1)-\ln(x+1)\] \[\ln(\frac{x-1}{x+1})\]

OpenStudy (turingtest):

I can do the integral just fine, but I'm unsure of the limit

OpenStudy (anonymous):

now i think it is ok, because as x goes to minus infinity you get \(\ln(1)=0\) i gotta run

OpenStudy (anonymous):

i don't think it's going to minus infinity, i think its going from minus infinity to minus 2

OpenStudy (anonymous):

i might be wrong

OpenStudy (turingtest):

it is... and I think satellite was in a hurry and messed up a little

OpenStudy (turingtest):

first the integral: partial fractions on the integrand\[{2\over x^2-1}={2\over(x+1)(x-1)}=\frac A{x+1}+\frac B{x-1}\]

OpenStudy (turingtest):

@sk84 do you know this technique^^ ?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

I think this diverges... integral is easy, limit... not so much

OpenStudy (turingtest):

so then you go ahead and get A and B while I think about the limit :)

OpenStudy (anonymous):

ok lol

OpenStudy (turingtest):

ok I think satellite is right about the limit

OpenStudy (turingtest):

let me know when you get A and B

OpenStudy (anonymous):

i think we have to do the lim of n->-2 from the left of ln (x-1) - ln (x+1)

OpenStudy (anonymous):

i dunno though

OpenStudy (zarkon):

you should write it as \[\ln\left(\frac{x-1}{x+1}\right)\] when you are evaluating the limits of integration

OpenStudy (zarkon):

just plug in \(x=-2\)

OpenStudy (turingtest):

...because as satellite said as \(x\to-\infty\) the expression becomes\[\ln1=0\]and we only have to evaluate at the other point ...right?

OpenStudy (turingtest):

*\(n\to-\infty\)

OpenStudy (zarkon):

\[\lim_{a\to-\infty}\int_{a}^{-2}{2dx\over x^2-1}\] \[=\lim_{a\to-\infty}\left.\ln\left(\frac{x-1}{x+1}\right)\right|_a^{-2}\] \[=\ln\left(\frac{-2-1}{-2+1}\right)-\lim_{a\to-\infty}\ln\left(\frac{a-1}{a+1}\right)\]

OpenStudy (turingtest):

Just so you know Zarkon, I had figured it out already. I was just elsewhere for a moment...

OpenStudy (zarkon):

good ;)

OpenStudy (zarkon):

\[=\ln(3)\]

OpenStudy (anonymous):

thanks guys! esp turing test and zarkon

OpenStudy (turingtest):

welcome!